Electronics and inhabitants of the International Space Station generate a significant amount of thermal energy that the station must get rid of. The only way that the station can exhaust thermal energy is by radiation, which it does using thin, 1.7 m -by-3.6 m panels that have a working temperature of about 6 ∘C. How much power is radiated from each panel? Assume that the panels are in the shade so that the absorbed radiation will be negligible. Take e=1.
Answer #1

here,

thickness of panels, t = 1.7 m by 3.6 m
area of panels, A = 1.7 * 3.6 = 6.12 m^2
working temp, T = 6 C = 273.15 + 6 = 279.15 K

Usng Stefan-Bolzmann law and assume an emissivity of 1.0

Radiated power, P :
P = (n)(e)(area)(sigma)(T^4)

where,
n, is no of sides
T is temp in kelvin
sigma is the Stefan-Boltzmann constant
A, is area

P = 2*1*6.12*5.67 * 10^-8*(279.15)^4
P = 4214.198 Watts or 4214.2 Watts (Rounded off)